OpenStudy (anonymous):

Can someone please help me? Find the exact value by using a half-angle identity. sin(7pi/8) its a practice question and im so confused.

OpenStudy (firejay5):

sin(7pi/8) = +_ ( sqrt( 1- cos(14pu/8)/2) = sqrt ( (1- cos7pi/4)/2 ) = sqrt ( 1-sqrt(3)/2 )/2 ) =sqrt (2 -sqrt(3))/2 It depends on the angel and assumption that you will take positive or negative

OpenStudy (anonymous):

@Firejay5 ...you literally just copied and pasted that from yahoo answers. i was hoping for an explanation but ok?????

OpenStudy (firejay5):

Well dude/girl did you want help you or not, that was up to you